# Adaptive resolution of fine scales in modes of microstructured optical fibers

### v1

@@ Abstract.
1. Introduction
@@
The eigenvalues of the Maxwell system for hybrid modes are squares of the physical propagation constant $\beta$ of the modes. Confinement loss of a fiber is a scalar multiple of the imaginary part of $\beta$. Typical values of $\beta$ have real parts of the order of $10^{6}$ and imaginary parts, $\text{Im}(\beta)$, of relevance that can range in orders of magnitude $10^{-6}$ to $10^{3}$, depending on the application. To get around this large variation in order of magnitudes, instead of a mere length scaling, we perform a nondimensionalization of the Maxwell system (see Section 2) akin to the process of going from Helmholtz to a Schrödinger problem. Then, instead of $\beta^{2}$, we compute a nondimensional $Z^{2}$ of unit order in the complex plane. Applying this approach to varied fibers, we find that an adaptive algorithm tailored to reduce the error in $Z^{2}$ also appears to get the confinement loss accurately and capture the above-mentioned fine-scale ripples. Whether a mathematically rigorous algorithm tailored specifically to reduce the error in loss can be found, and if so, whether it will do better, is an issue for further research.
The basis for DWR-based adaptive algorithms is developed in terms of optimal control in the review [1]. Roughly speaking, their idea is that the error in a quantity of interest can be estimated by an element-wise residual (which treats local error contributions), weighted by the solution of a dual problem (which takes global error transport into account). DWR algorithms specific for eigenvalues have been outlined for elliptic problems in [12, 26]. We bring in an additional technique, the use of a regular decomposition, that is useful to handle our Maxwell example. The essential theory is summarized in a single self-contained theorem and proof in Section 3 giving the error estimator expression, after which we describe our adaptive algorithm. Later sections clearly reveal the efficacy of the algorithm for computing leaky modes of complex fibers. In particular, Section 4 is devoted to verifying the correctness of the methodology using a Bragg fiber for which modes in exact closed form are available.
@@
Figure 1. Transverse geometries of the microstructured fibers (not to scale) studied in this work. White regions indicate air and shades of blue indicate dielectric materials with higher refractive indices.
@@
2. The eigenproblem for leaky Maxwell modes
@@
for some propagation constant $\beta$ and vector fields $E,H$ to be determined. By using unit vectors $e_{0},e_{1},e_{2}\in\mathbb{R}^{3}$ in the coordinate directions, the transverse part of a vector field $F=F(x_{0},x_{1},x_{2})$ is its projection into the span of $e_{0}$ and $e_{1},$ denoted by $F_{\tau}$, so that $F=F_{\tau}+F_{2}e_{2}$ where $F_{2}$ is the longitudinal component of $F$. Let $R$ denote the 90 degree rotation operator that maps any transverse field $F_{\tau}=F_{0}e_{0}+F_{1}e_{1}$ to $RF_{\tau}=F_{1}e_{0}-F_{0}e_{1}.$ Define the divergence and the scalar-valued curl for transverse fields by $\operatorname{div}F_{\tau}=\partial_{0}F_{0}+\partial_{1}F_{1}$ and $\operatorname{curl}(F_{\tau})=\operatorname{div}(RF_{\tau}).$ The rotated gradient of a scalar field $\phi$ is defined by $\operatorname{rot}\phi=R\ \operatorname{grad}(\phi).$ Next, we substitute (2) into (1), decompose $E$ and $H$ into transverse and longitudinal components, and use the identity ×$\nabla\times\left(u\,e^{\imath\beta x_{2}}\right)=\left(e^{\imath\beta x_{2}}\operatorname{curl}u_{\tau}\right)e_{2}+e^{\imath\beta x_{2}}\left(\operatorname{rot}u_{2}-\imath\beta Ru_{\tau}\right).$ Then, eliminating $H_{2}$ and $H_{\tau}$ and simplifying (see e.g., [30, Appendix A.1] for more details), the Maxwell system reduces to a system coupling the transverse electric field $E_{\tau}$ and a scaled longitudinal component of the field $\varphi=\imath\beta E_{2}$, which reads as follows:
@@
The physical dimensions of the fiber’s transverse cross-section are in micrometers, while practical propagation constants are usually of the order of $10^{6}$, so a nondimensionalization is useful. Let $\varepsilon_{0}$ and $\mu_{0}$ be the electric permittivity and magnetic permeability of the vacuum, respectively. Let $k^{2}=\omega^{2}\varepsilon_{0}\mu_{0}$. Define the transverse and longitudinal refractive indices by $n_{\tau}^{2}=\varepsilon_{\tau}\mu_{0}/\varepsilon_{0}\mu$ and $n_{2}^{2}=\varepsilon_{2}\mu_{0}/\varepsilon_{0}\mu$. Now, using a characteristic length scale $L$ for the fiber cross-section, we rescale $(x_{0},x_{1})\mapsto\left(x_{0}/L,x_{1}/L\right)$ and consider all the unknown functions in the rescaled coordinates. Outside some finite radius (of the same order as $L$) in the transverse plane we assume that the refractive index $n(x_{0},x_{1})$ is isotropic and equal to a constant $n_{0}$. Using an index well defined by $V(x_{0},x_{1})=L^{2}k^{2}(n_{0}^{2}-n_{\tau}(x_{0},x_{1})^{2})$ in the rescaled nondimensional coordinates, the system (3) leads us [30, Appendix A.2] to the following non-dimensional equations:
@@
The system (4) must be supplemented with boundary conditions. Before implementing the outgoing conditions of leaky modes, it is useful to consider the case where all fields have decayed to some negligible magnitude (as in a guided mode) at the boundary of a large enough disk $\Omega$ in the $x_{0}x_{1}$-plane, sufficiently far from the compact support of the index well $V$. Multiplying the first equation of (4) by a vector test function $F$ and the second by a scalar test function $\psi$, integrating by parts, and using $E_{\tau}=0$ and $\varphi=0$ on $\partial\Omega$, we obtain the following weak formulation. Find $E_{\tau}\in\mathring{H}(\operatorname{curl},\Omega)$ and $\varphi\in\mathring{H}^{1}(\Omega)$, satisfying
(5a) $(\operatorname{curl}E_{\tau},\operatorname{curl}F)_{\Omega}+(VE_{\tau},F)_{\Omega}+(\operatorname{grad}\varphi,F)_{\Omega}$ $=Z^{2}(E_{\tau},F)_{\Omega},$
(5b) $(n_{2}^{2}\varphi,\psi)_{\Omega}-(n_{\tau}^{2}E_{\tau},\operatorname{grad}\psi)_{\Omega}$ $=0,$
@@
for all $F\in\mathring{H}(\operatorname{curl},\Omega)$ and all $\psi\in\mathring{H}^{1}(\Omega)$. Here $(\cdot,\cdot)_{\Omega}$ denotes the (complex) inner product of $L^{2}(\Omega)$ or its Cartesian products. Here $\mathring{H}(\operatorname{curl},\Omega)$ denotes the space of square-integrable vector fields whose $\operatorname{curl}$ is also square integrable and whose tangential component along $\partial\Omega$ vanishes and $\mathring{H}^{1}(\Omega)$ denotes square-integrable scalar functions which vanish on $\partial\Omega$ all of whose first-order derivatives are also square integrable on $\Omega$. Prior works featuring similar weak formulations include [16, 17, 30, 31].
@@
Leaky modes satisfy (4) and also satisfy the condition at infinity that the mode must be outgoing. A convenient method to impose this condition is through a perfectly matched layer (PML) [2] which transforms coordinates so that the solution of (4), while unaltered in a bounded region where $V$ is inhomogeneous, becomes exponentially decaying outside. The exponential decay allows us to truncate the infinite domain $\mathbb{R}^{2}$ in (4) to a bounded computational domain $\Omega$ as in (5). To view PML as a complex coordinate transformation [20, 5], we let $\tilde{x}:=\Phi(x)$ where $\Phi(x)=x\,\eta(r)/r$ for some complex-valued function $\eta(r)$ to be specified later. Here $x=[x_{0},x_{1}]^{\mathsf{T}}$, $\tilde{x}=[\tilde{x}_{0},\tilde{x}_{1}]^{\mathsf{T}}$, and $r=\sqrt{x_{0}^{2}+x_{1}^{2}}$. Put $x_{\bot}=[x_{1},-x_{0}]^{\mathsf{T}}$. Consider the $2\times 2$ Jacobian matrix whose entries are $J_{ij}:=\partial\tilde{x}_{i}/\partial x_{j}$. Simple computations give that with $\zeta(r)=\eta(r)/r$, we have
@@
By placing a tilde over a differential operator, we indicate that the derivative is being taken with respect to $\tilde{x}$ rather than $x$. Calculations using the chain rule exhibits that the matrix $D\Psi$ whose $(i,j)$th entry equals $\partial_{j}\Psi_{i}$ satisfies ~$\tilde{\partial}_{j}\tilde{\Psi}_{i}=[J^{-\mathsf{T}}\ D\Psi\ J^{-1}]_{ij}+\sum_{l}\Psi_{l}\,\partial^{2}x_{l}/\partial\tilde{x}_{j}\partial\tilde{x}_{i}.$ Relating its skew symmetric part to the curl and applying (6), we obtain
@@
(9a) $(\tilde{\operatorname{curl}}\tilde{E}_{\tau},\tilde{\operatorname{curl}}\tilde{F})_{\tilde{\Omega}}+(V\tilde{E}_{\tau},\tilde{F})_{\tilde{\Omega}}+(\operatorname{grad}\tilde{\varphi},\tilde{F})_{\tilde{\Omega}}$ $=Z^{2}(\tilde{E}_{\tau},\tilde{F})_{\tilde{\Omega}},$
(9b) $(n_{2}^{2}\tilde{\varphi},\tilde{\psi})_{\tilde{\Omega}}-(n_{\tau}^{2}\tilde{E}_{\tau},\operatorname{grad}\tilde{\psi})_{\tilde{\Omega}}$ $=0.$
@@
$(\tilde{\operatorname{curl}}\tilde{E},\tilde{\operatorname{curl}}\tilde{F})_{\tilde{\Omega}}=(\kappa\operatorname{curl}E,\operatorname{curl}F)_{\Omega}$
@@
(13b) $b(E_{\tau},\psi)$ $=\left(n_{\tau}^{2}\gamma E_{\tau},\operatorname{grad}\psi\right)_{\Omega},$ $c(\phi,F)=\left(\gamma\operatorname{grad}\phi,F\right)_{\Omega},\qquad$
(13c) $d(\phi,\psi)$ $=-\left(n_{2}^{2}{\zeta}\eta^{\prime}\phi,\psi\right)_{\Omega},$ $m(E_{\tau},F)=\left(\gamma E_{\tau},F\right)_{\Omega}.$
The system (12) completes the description of the weak formulation for computing leaky modes, except for a prescription of $\eta(r)$.
@@
We conclude this section by describing the finite element discretization of (12) that we shall employ. First, mesh $\Omega$ by a geometrically conforming finite element mesh of triangles, and denote the mesh by $\Omega_{h}$. On a triangle $K$, let $P_{p}(K)$ denote the space of polynomials in two variables of degree at most $p$. The degree $p$ and the maximal element diameter $h=\max_{K\in\Omega_{h}}\mathrm{diam}(K)$ determine the richness of the discretization. The Lagrange finite element space on $\Omega_{h}$, for any $p\geq 0$ is defined by $V_{h}=\{\psi:\psi|_{K}\in P_{p+1}(K)$ for all $K\in\Omega_{h},$ $\psi$ is continuous}.}\}.} . The Nédélec finite element space [20] is the set of all vector fields $F$ on $\Omega$ whose tangential components are continuous across element interfaces and whose restrictions to each element $K\in\Omega_{h}$ is a polynomial of the form $F|_{K}\in P_{p}(K)^{2}+\left[\begin{smallmatrix}\phantom{-}x_{1}\\ -x_{0}\end{smallmatrix}\right]P_{p}(K).$ We use curved triangles next to curved material interfaces, in which case, as usual, the polynomial space within such an element is replaced by the pullback of the above-indicated polynomial spaces from a reference triangle. The Galerkin discretization using these spaces give the discrete eigenproblem that solves for $E_{\tau,h}\in N_{h}$ and $\varphi_{h}\in V_{h}$ such that
@@
3. The error estimator for eigenvalues

### v2

@@ Abstract
keywords:
optical fibers, finite element adaptivity, mode solving, vector hybrid leaky modes, mode confinement losses, microstructured fibers
{MSCcodes}
65N30, 78A10
1 Introduction
@@
The eigenvalues of the Maxwell system for hybrid modes are squares of the physical propagation constant $\beta$ of the modes. Confinement loss of a fiber is a scalar multiple of the imaginary part of $\beta$. Typical values of $\beta$ have real parts of the order of $10^{6}$ and imaginary parts, $\text{Im}(\beta)$, that can range in orders of magnitude $10^{-6}$ to $10^{3}$, depending on the fiber and operating wavelength. To get around this large variation in order of magnitudes, instead of a mere length scaling, we perform a nondimensionalization of the Maxwell system (see Section 2) akin to the process of going from Helmholtz to a Schrödinger problem. Then, instead of $\beta^{2}$, we compute a nondimensional $Z^{2}$, of unit order of magnitude, in the complex plane. Applying this approach to varied fibers, we find that an adaptive algorithm tailored to reduce the error in $Z^{2}$ also appears to get the confinement loss accurately and capture the above-mentioned fine-scale ripples. Whether a mathematically rigorous algorithm tailored specifically to reduce the error in loss can be found, and if so, whether it will do better, is an issue for further research.
The basis for DWR-based adaptive algorithms is developed in terms of optimal control in the review [1]. Roughly speaking, their idea is that the error in a quantity of interest can be estimated by an element-wise residual (which treats local error contributions), weighted by the solution of a dual problem (which takes global error transport into account). DWR algorithms specific for eigenvalues have been outlined for elliptic problems in [12, 26]. We choose the DWR error estimation technique because our eigenproblem is nonselfadjoint and because our eigensolver readily gives the dual solution (left eigenvector) in addition to the solution (the eigenvalue and the right eigenvector). Not incurring extra costs for solving the dual problem in this case makes the DWR technique an efficient choice. To apply the DWR methodology to our Maxwell eigenvalue application, we bring in an additional technique, namely the use of a regular decomposition. The resulting theory is summarized in a single self-contained theorem and proof in Section 3 giving the error estimator expression, after which we describe our adaptive algorithm. Later sections clearly reveal the efficacy of the algorithm for computing leaky modes of complex fibers. In particular, Section 4 is devoted to verifying the correctness of the methodology using a Bragg fiber for which mode solutions in exact closed form are available.
@@
Figure 1: Transverse geometries of the microstructured fibers (not to scale) studied in this work. White regions indicate air and shades of blue indicate dielectric materials with higher refractive indices.
@@
2 The eigenproblem for leaky Maxwell modes
@@
for some propagation constant $\beta$ and vector fields $E,H$ to be determined. By using unit vectors $e_{0},e_{1},e_{2}\in\mathbb{R}^{3}$ in the coordinate directions, the transverse part of a vector field $F=F(x_{0},x_{1},x_{2})$ is its projection into the span of $e_{0}$ and $e_{1},$ denoted by $F_{\tau}$, so that $F=F_{\tau}+F_{2}e_{2}$ where $F_{2}$ is the longitudinal component of $F$. Let $R$ denote the 90 degree rotation operator that maps any transverse field $F_{\tau}=F_{0}e_{0}+F_{1}e_{1}$ to $RF_{\tau}=F_{1}e_{0}-F_{0}e_{1}$. Define the divergence and the scalar-valued curl for transverse fields by $\operatorname{div}F_{\tau}=\partial_{0}F_{0}+\partial_{1}F_{1}$ and $\operatorname{curl}(F_{\tau})=\operatorname{div}(RF_{\tau}).$ The rotated gradient of a scalar field $\phi$ is defined by $\operatorname{rot}\phi=R\ \operatorname{grad}(\phi)$. Next, we substitute (2) into (1), decompose $E$ and $H$ into transverse and longitudinal components, and use the identity ×$\nabla\times\left(u\,e^{\imath\beta x_{2}}\right)=\left(e^{\imath\beta x_{2}}\operatorname{curl}u_{\tau}\right)e_{2}+e^{\imath\beta x_{2}}\left(\operatorname{rot}u_{2}-\imath\beta Ru_{\tau}\right)$. Then, eliminating $H_{2}$ and $H_{\tau}$ and simplifying (see e.g., [31, Appendix A.1] for more details), the Maxwell system reduces to a system coupling the transverse electric field $E_{\tau}$ and a scaled longitudinal component of the field $\varphi=\imath\beta E_{2}$, which reads as follows:
@@
The physical dimensions of the fiber’s transverse cross-section are in micrometers, while practical propagation constants are usually of the order of $10^{6}$, so a nondimensionalization is useful. Let $\varepsilon_{0}$ and $\mu_{0}$ be the electric permittivity and magnetic permeability of the vacuum, respectively. Let $k^{2}=\omega^{2}\varepsilon_{0}\mu_{0}$. Define the transverse and longitudinal refractive indices by $n_{\tau}^{2}=\varepsilon_{\tau}\mu_{0}/\varepsilon_{0}\mu$ and $n_{2}^{2}=\varepsilon_{2}\mu_{0}/\varepsilon_{0}\mu$. Now, using a characteristic length scale $L$ for the fiber cross-section, we rescale $(x_{0},x_{1})\mapsto\left(x_{0}/L,x_{1}/L\right)$ and consider all the unknown functions in the rescaled coordinates. Outside some finite radius (of the same order as $L$) in the transverse plane we assume that the refractive index $n(x_{0},x_{1})$ is isotropic and equal to a constant $n_{0}$. Using an index well defined by $V(x_{0},x_{1})=L^{2}k^{2}(n_{0}^{2}-n_{\tau}(x_{0},x_{1})^{2})$ in the rescaled nondimensional coordinates, the system (3) leads us [31, Appendix A.2] to the following non-dimensional equations:
@@
The system (4) must be supplemented with boundary conditions. We shall use the outgoing condition at infinity, described below. Before discussing it however, consider the case where all fields have decayed to some negligible magnitude (as in a guided mode) at the boundary of a large enough disk $\Omega$ in the $x_{0}x_{1}$-plane, sufficiently far from the compact support of the index well $V$. Multiplying the first equation of (4) by a vector test function $F$ and the second by a scalar test function $\psi$, integrating by parts, and using $E_{\tau}=0$ and $\varphi=0$ on $\partial\Omega$, we obtain the following weak formulation. Find $E_{\tau}\in\mathring{H}(\operatorname{curl},\Omega)$ and $\varphi\in\mathring{H}^{1}(\Omega)$, satisfying
(5a) $\int_{\Omega}(\operatorname{curl}E_{\tau})(\operatorname{curl}F)+\int_{\Omega}VE_{\tau}\cdot F+\int_{\Omega}(\operatorname{grad}\varphi)\cdot F$ $=Z^{2}\int_{\Omega}E_{\tau}\cdot F,$
(5b) $\int_{\Omega}n_{2}^{2}\varphi\psi-\int_{\Omega}n_{\tau}^{2}E_{\tau}\cdot\operatorname{grad}\psi$ $=0,$
@@
for all $F\in\mathring{H}(\operatorname{curl},\Omega)$ and all $\psi\in\mathring{H}^{1}(\Omega)$. Here $\mathring{H}(\operatorname{curl},\Omega)$ denotes the space of square-integrable vector fields whose $\operatorname{curl}$ is also square integrable and whose tangential component along $\partial\Omega$ vanishes and $\mathring{H}^{1}(\Omega)$ denotes square-integrable scalar functions which vanish on $\partial\Omega$ all of whose first-order derivatives are also square integrable on $\Omega$. Prior works featuring similar weak formulations include [16, 17, 31, 32].
@@
Leaky modes satisfy (4) and also satisfy the condition at infinity that the mode must be outgoing. To explain this, first note that the fiber’s translational symmetry in $x_{2}$-direction and the absence of inhomogeneities in $r^{2}\equiv x_{0}^{2}+x_{1}^{2}>R^{2}$ for some finite cylindrical radius $R$, implies that we can use separation of variables to obtain a series expansion of solution components in terms of solutions of the Bessel equation in $r$ and sines and cosines in the angular variable $\theta$. From among all solutions to the Bessel equation, when $Z$ is a positive real number, the outgoing requirement selects Hankel functions of the first kind, $H_{\ell}^{(1)}$. Hence the radial variation of all terms in the solution representation are of the form $H_{\ell}^{(1)}(Zr)$ as $r\to\infty$ (see e.g., [5] or [10, eq. (23)]). For complex values of $Z$, the outgoing condition requires that the solution is expressible in terms of the analytic continuations into $\mathbb{C}$, from the positive real axis, of the same Hankel functions $H_{\ell}^{(1)}(Zr)$. A practically convenient method to impose this condition is through a perfectly matched layer (PML) [2] which transforms coordinates so that the solution of (4), while unaltered in a bounded region where $V$ is inhomogeneous, becomes exponentially decaying outside. The exponential decay allows us to truncate the infinite domain $\mathbb{R}^{2}$ in (4) to a bounded computational domain $\Omega$ as in (5). To view PML as a complex coordinate transformation [20, 5], we let $\tilde{x}:=\Phi(x)$ where $\Phi(x)=x\,\eta(r)/r$ for some complex-valued function $\eta(r)$ to be specified later. Here $x=[x_{0},x_{1}]^{\mathsf{T}}$, $\tilde{x}=[\tilde{x}_{0},\tilde{x}_{1}]^{\mathsf{T}}$, and $r=\sqrt{x_{0}^{2}+x_{1}^{2}}$. Put $x_{\bot}=[x_{1},-x_{0}]^{\mathsf{T}}$. Consider the $2\times 2$ Jacobian matrix whose entries are $J_{ij}:=\partial\tilde{x}_{i}/\partial x_{j}$. Simple computations give that with $\zeta(r)=\eta(r)/r$, we have
@@
By placing a tilde over a differential operator, we indicate that the derivative is being taken with respect to $\tilde{x}$ rather than $x$. Calculations using the chain rule shows that the matrix $D\Psi$ whose $(i,j)$th entry equals $\partial_{j}\Psi_{i}$ satisfies ~$\tilde{\partial}_{j}\tilde{\Psi}_{i}=[J^{-\mathsf{T}}\ D\Psi\ J^{-1}]_{ij}+\sum_{l}\Psi_{l}\,\partial^{2}x_{l}/\partial\tilde{x}_{j}\partial\tilde{x}_{i}$. Relating its skew symmetric part to the curl and applying (6), we obtain
@@
(9a) $\int_{\tilde{\Omega}}(\tilde{\operatorname{curl}}\tilde{E}_{\tau})(\tilde{\operatorname{curl}}\tilde{F})+\int_{\tilde{\Omega}}V\tilde{E}_{\tau}\cdot\tilde{F}+\int_{\tilde{\Omega}}(\operatorname{grad}\tilde{\varphi})\cdot\tilde{F})$ $=Z^{2}\int_{\tilde{\Omega}}\tilde{E}_{\tau}\cdot\tilde{F},$
(9b) $\int_{\tilde{\Omega}}n_{2}^{2}\tilde{\varphi}\tilde{\psi}-\int_{\tilde{\Omega}}n_{\tau}^{2}\tilde{E}_{\tau}\cdot\operatorname{grad}\tilde{\psi}$ $=0.$
@@
$\int_{\tilde{\Omega}}(\tilde{\operatorname{curl}}\tilde{E}_{\tau})(\tilde{\operatorname{curl}}\tilde{F})=\int_{\tilde{\Omega}}\kappa\,\operatorname{curl}E_{\tau}\,\operatorname{curl}F,$
@@
(13b) $b(E_{\tau},\psi)$ $=\left(n_{\tau}^{2}\gamma E_{\tau},\operatorname{grad}\psi\right)_{\Omega},\qquad c(\phi,F)=\left(\gamma\operatorname{grad}\phi,F\right)_{\Omega},\qquad$
(13c) $d(\phi,\psi)$ $=-\left(n_{2}^{2}{\zeta}\eta^{\prime}\phi,\psi\right)_{\Omega},\qquad\qquad m(E_{\tau},F)=\left(\gamma E_{\tau},F\right)_{\Omega}.$
Here $(\cdot,\cdot)_{\Omega}$ denotes the (complex) inner product of $L^{2}(\Omega)$ or its Cartesian products and we have used the conjugates of test functions in (9). The system (12) completes the description of the weak formulation for computing leaky modes, except for a prescription of $\eta(r)$.
@@
We conclude this section by describing the finite element discretization of (12) that we shall employ. First, mesh $\Omega$ by a geometrically conforming finite element mesh of triangles, and denote the mesh by $\Omega_{h}$. On a triangle $K$, let $P_{p}(K)$ denote the space of polynomials in two variables of degree at most $p$. The degree $p$ and the maximal element diameter $h=\max_{K\in\Omega_{h}}\mathrm{diam}(K)$ determine the richness of the discretization. The Lagrange finite element space on $\Omega_{h}$, for any $p\geq 0$ is defined by $V_{h}=\{\psi:\psi|_{K}\in P_{p+1}(K)$ for all $K\in\Omega_{h},$ $\psi$ is continuous}}\}}. The Nédélec finite element space [20] is the set of all vector fields $F$ on $\Omega$ whose tangential components are continuous across element interfaces and whose restrictions to each element $K\in\Omega_{h}$ is a polynomial of the form $F|_{K}\in P_{p}(K)^{2}+\left[\begin{smallmatrix}\phantom{-}x_{1}\\ -x_{0}\end{smallmatrix}\right]P_{p}(K)$. We use curved triangles next to curved material interfaces, in which case, as usual, the polynomial space within such an element is replaced by the pullback of the above-indicated polynomial spaces from a reference triangle. The Galerkin discretization using these spaces give the discrete eigenproblem that solves for $E_{\tau,h}\in N_{h}$ and $\varphi_{h}\in V_{h}$ such that
@@
3 The error estimator for eigenvalues